OpenStudy (anonymous):

A tropical cream is sold in a 26-gram container. The average amount of tropical cream used for application is 2/5 grams. How many applications can be made with the container?

OpenStudy (amistre64):

26/(2/5) 26(5)/2 = 13(5) = 50 + 15 = 65

OpenStudy (amistre64):

chaeck: 65/5 = 13(2) = 26

OpenStudy (anonymous):

i got it, thanx

OpenStudy (amistre64):

youre welcome :)